Subshilava:

PLEASE HELP!! Look at the square ABCD below. A(- 2, 0) B(0, 6) C(6,4) and D(4, - 2) What is its perimeter? 8sqrt(40) 8sqrt(10) 4sqrt(80) 4sqrt(10)

toga:

Where is the square

toga:

please send a pic of it

Subshilava:

1 attachment
Subshilava:

@toga wrote:
please send a pic of it
I did

toga:

ok i will look at it

Subshilava:

Thank you :)

toga:

what grade are you in

Subshilava:

Sophomore

toga:

oh that is why I don't remember learning this I can try to help but I might be wrong

Subshilava:

Ok your fine I learned some of this last year but I don’t really remember any of it

toga:

To find the perimeter of the square, we need to find the distance between its four sides, which are the line segments connecting points A to B, B to C, C to D, and D to A. Using the distance formula: distance = sqrt((x2 - x1)^2 + (y2 - y1)^2) We can find the distance between points A and B as: AB = sqrt((0 - (-2))^2 + (6 - 0)^2) AB = sqrt(4^2 + 6^2) AB = sqrt(52) AB = 2sqrt(13) Similarly, we can find the distance between points B and C as: BC = sqrt((6 - 0)^2 + (4 - 6)^2) BC = sqrt(6^2 + 2^2) BC = 2sqrt(10) We can find the distance between points C and D as: CD = sqrt((4 - 6)^2 + (-2 - 4)^2) CD = sqrt((-2)^2 + (-6)^2) CD = 2sqrt(10) Finally, we can find the distance between points D and A as: DA = sqrt((-2 - (-2))^2 + (-2 - 0)^2) DA = sqrt(0^2 + (-2)^2) DA = 2 Therefore, the perimeter of the square is: AB + BC + CD + DA = 2sqrt(13) + 2sqrt(10) + 2sqrt(10) + 2 Perimeter = 2sqrt(13) + 4sqrt(10) + 2 I hope this helps!